# Sub-nanometer 3D morphometric precision of polarisation-resolved wide-field optical extinction microscopy determines the roundness of individual gold nanospheres

### v1

@@ IV Results and Discussion
This work demonstrates the enhanced capabilities of the new ON including radial polarisation in determining individual NP morphometries in 3D. We maintain the naming and definitions of variables and quantities used in our previous publications,[43, 44] and introduce them in the text where needed. All analysis assumes that the refractive index of the NP environment in the samples is $n_{\mathrm{m}}=1.52$.
@@ IV.1 Statistical Distributions of $\sigma_{\Lambda}$ and $\alpha_{\Lambda}$
We first introduce relevant quantities used in the statistical description of the optical measurements. For each NP, Eq. 9 is fitted to the $\sigma_{\Lambda}(\gamma_{\mathrm{P}})$ measured for linear polarisation, determining its parameters $\sigma$, $\alpha$, and $\gamma$. Their standard errors $\hat{\sigma}$, $\hat{\alpha}$, and $\hat{\gamma}$ are determined using a Monte Carlo simulation where we add random Gaussian noise to the measured data, matching the experimentally determined noise, as detailed in Ref. 43 and Sec. LABEL:SM-sec:optnoise.
Given a set of $N$ NPs, numbered by the index, $i$, we identify the fit parameters using the subscript $\Lambda,i$, and define the sample-wide mean of $\sigma_{\Lambda,\,i}$ to be $\bar{\sigma}_{\Lambda}=\frac{1}{N}\sum_{i=1}^{N}\sigma_{\Lambda,\,i}$, and the associated standard deviation $\hat{\sigma}_{\Lambda,\mathrm{m}}$. Equivalent quantities are defined for the other two parameters. Correcting for the measurement noise, the standard deviation $\hat{\sigma}_{\Lambda}=\sqrt{\hat{\sigma}_{\Lambda,\mathrm{m}}^{2}-\frac{1}{N}\sum_{i=1}^{N}\hat{\sigma}_{\mathrm{\Lambda},i}^{2}}$ represents the distribution of NP cross-sections originating from their sizes and shapes only. For the relative amplitude parameter the measurement error $\hat{\alpha}_{\Lambda,i}$ varies strongly depending on the parameter values, so that we refrain from such a correction and directly use $\hat{\alpha}_{\Lambda}=\hat{\alpha}_{\Lambda,\,\mathrm{m}}$.
@@
Fig. 2 presents $\sigma_{\Lambda}$ and $\alpha_{\Lambda}$ for measurements of the $N=51$ nominally D=30D=30\,nm UGNS for all $\Lambda$ of the experiment. To mitigate small differences in focus between the different fields of view, $\sigma_{\mathrm{ext}}$ was determined using the larger radius $r_{\mathrm{h}}$ (see Sec. LABEL:SM-sec:sigcomp for more details). One must correct $\sigma_{\mathrm{ext}}$ obtained with a finite $r_{\mathrm{i}}$ to compensate for the signal outside of the integration area. The correction factor to $\sigma_{\mathrm{ext}}$ was determined to be $c_{\mathrm{l}}=1.12$, $c_{\mathrm{s}}=1.06$, and $c_{\mathrm{h}}=1.02$, for $r_{\mathrm{l}}$, $r_{\mathrm{s}}$, and $r_{\mathrm{h}}$, respectively. In Fig. 2 we show $\sigma_{\mathrm{ext}}$ as measured and the modelled extinction divided by $c_{\mathrm{h}}$. Narrow distributions of $\sigma_{\Lambda}$ are observed for all $\Lambda$. Small in-plane asymmetries $\alpha_{\Lambda}<0.2$ are found for about 95% of the NPs at $\Lambda={\color[rgb]{1,0.6,0}600}$ and below, and for about 80% of NPs at $\Lambda={\color[rgb]{1,0,0}650}$. At $\Lambda={\color[rgb]{0.7,0,0}700}$ instead, $\sigma_{\mathrm{ext}}$ is only a few times larger than the noise, which results in a wide distribution of $\alpha_{\Lambda}$ estimated by the shaded areas. Hence, the sample exhibits a low optical asymmetry, with some 10% outliers.
@@
Looking at the simulations in the inset of Fig. 2, we note that $\Lambda={\color[rgb]{0,0.9,0}550}$ covers the higher wavelength side of the LSPR of a spherical gold NP, suggesting that accurately modelling the width of the channels will be relevant for the morphology retrieval,. With increasing NP asymmetry, the LSPR polarised along the longer axis (the longitudinal mode) red-shifts, increasing its overlap with the $\color[rgb]{0,0.9,0}550$} channel, and thus providing a sensitive probe of asymmetries. We find $\bar{\alpha}_{{\color[rgb]{0,0.9,0}550}}=0.037\,\pm\,0.039$, with some 80% of NPs showing $\alpha_{{\color[rgb]{0,0.9,0}550}}<0.054$, and the rest clearly asymmetric with $\alpha_{{\color[rgb]{0,0.9,0}550}}$ up to 0.194. For asymmetric NPs, the LSPR will move into the increasingly red channels (see simulations in Ref. 43 Fig. 2), increasing their $\sigma_{\Lambda}$ and $\alpha_{\Lambda}$. Looking at the exemplary TEM image, some 10-20% of significantly non-spherical NPs can be expected. Their red-shift can be significant, contributing to the values of $\alpha_{{\color[rgb]{1,0,0}650}}$ and $\alpha_{{\color[rgb]{0.7,0,0}700}}$ well above the noise. We also note that the $\Lambda=\{{\color[rgb]{0,0,1}450},\,{\color[rgb]{0,0.92,0.92}500}\}$ channels are rather insensitive to asymmetry, due to the dominant absorptive part of the permittivity, $\Im\{\epsilon_{\mathrm{NP}}\}\approx 5$, with $\Re\{\epsilon_{\mathrm{NP}}\}$ similar to $\epsilon_{\mathrm{m}}$, suppressing shape-dependent effects, so that $\sigma_{{\color[rgb]{0,0,1}450}}$ and $\sigma_{{\color[rgb]{0,0.92,0.92}500}}$ are good reporters of the NP volume.
@@
Let us now discuss $\sigma_{\mathrm{ext}}$ measured using the radial polariser, $\sigma_{\Lambda}(\gamma_{\mathrm{R}})$, see vertical bars in Fig. 2, which we compare to the fitted $\sigma_{\Lambda}$ for linear polarisation shown as crosses. We recall that the electric field at the sample has both in-plane and out-of-plane components, and we consider[35, 52] the relative strengths of the components along the $x$, $y$, and $z$ axes, respectively as $E=[E_{\mathrm{x}},\,E_{\mathrm{y}},\,E_{\mathrm{z}}]$. The intensity in the back focal plane (BFP) of the condenser varies spatially, with a maximum at the center, determined by the lamp, diffuser and lenses in the illumination path, as characterised in our previous work.[35] Using the known spatial dependence, we can calculate the intensity components at the sample polarised along the $x$, $y$, and $z$ axes, normalized to have unity sum. For a linear polariser along $x$ in the condenser BFP, we find of $I_{\mathrm{lin}}=[0.826,\,0.007,\,0.167]$, corresponding to field components $E_{\mathrm{lin}}=[0.909,\,0.084,\,0.409]$ given by the square root of the intensity components. For radially polarised input to the BFP, we have $I_{\mathrm{rad}}=[0.333,0.333,0.334]$ corresponding to $E_{\mathrm{rad}}=[0.577,\,0.577,\,0.578]$, showing that the $z$ component is increased in the radially polarised case.
@@
For a spherical NP, the polarisation direction should not affect the measured $\sigma_{\mathrm{ext}}$, and thus we expect the difference $\delta_{\sigma}=\sigma_{\Lambda}(\gamma_{\mathrm{R}})-\sigma_{\Lambda}$ to be negligible. The histograms of $\delta_{\sigma}$ are shown in Fig. 2b, and we find for channels with large cross-sections $\Lambda=\{{\color[rgb]{0,0,1}450},\,{\color[rgb]{0,0.92,0.92}500},\,{\color[rgb]{0,0.9,0}550}\}$ a shift towards $\delta_{\sigma}>0$ of relative magnitude $\delta_{\sigma}/\sigma_{\Lambda}$ below 3%, while for the other channels $\delta_{\sigma}$ is dominated by noise, shown as gray bars. To explain this small shift we recall the “long shadow” effect[55] which arises from the illumination covering an oblique angular range while referencing to the intensity transmitted through the imaged sample plane. The $\mathrm{NA}$ of the condenser and the intensity distribution in its BFP provide the weighting of the illumination angles at the sample, from which we calculate the long-shadow factor used. However, the transmission of the condenser and objective for oblique angles is polarisation-dependent, due to the reflections on the lens surfaces, which are generally lower for p polarisation compared to s polarisation. Radial polarisation is purely p polarised, while linear polarisation contains equal parts of s and p polarised light. Considering that the long-shadow factor is increasing with incidence angle,[55] the higher transmission at higher angles for radial polarisation creates an illumination with a higher long-shadow factor. We explore in Fig. LABEL:SM-fig:multi the shape retrieval when compensating the shift by reducing the measured $\sigma_{\Lambda}(\gamma_{\mathrm{R}})$ by a factor $0.97$, showing a slightly improved shape retrieval. One could also speculate that a preferential orientation of the NPs on the surface creates the difference. However, we would expect in such a case that NPs preferentially lie flat on the surface, providing a larger binding surface, which would lead to a higher cross-section for in-plane fields, opposite to what is observed.
@@
Figure 3: $\bar{S}$ and selected statistics of $S$ for $N=51$ UGNS of nominal D=30D=30\,nm for varying surface damping parameter $g$ and different $\epsilon$ datasets as indicated by color. (a) Top: $\bar{S}$ versus $g$; bottom: $M\{\mathrm{AR}\}$, the median of the fitted average aspect ratio, AR $\sqrt{bc}/a$. (b–j) Histograms of $S$, with (b–d) using the OL dataset with $g=\{0,\,1.8,\,3.6\}$, (e–g) JC dataset with $g=\{0,\,0.3,\,1.8\}$, (h–j) MP dataset with $g=\{0,\,1.8,\,3.6\}$. The values of $g$ shown are at $g=0$, the minimum of $\bar{S}(g)$ and a larger $g$.
@@ IV.2 Morphometric analysis
where we dropped the index $i$ for brevity. $n$ is the number of measurements in the sum, which includes the radially polarised measurement $\gamma_{\mathrm{P}}=\gamma_{\mathrm{R}}$, and $\hat{\sigma}_{\Lambda}(\gamma_{\mathrm{P}})$ is the measurement noise associated with $\sigma_{\Lambda}$($\gamma_{\mathrm{P}}$). The factor $\eta$ is the fraction of scattered light which is not collected by the objective and thus contributes to the measured extinction. For our experimental configuration we calculated [54] $\eta=0.638$.
@@
with $M\{.\}$ indicating the median over the $N$ NPs analyzed. Since $S$ is normalized by $\hat{\sigma}_{\mathrm{ext}}(\gamma_{\mathrm{P}})$ in Eq. 11, we would expect $S=1$ for the correct model, without taking into account the number of fit parameters and the number of data points to be fitted per NP. Considering that we use $p=6$ free parameters to fit the $n=30$ measurements (four linear and one radial polarisation at 6 wavelengths), this expectation is reduced by the number of degrees of freedom npn-p to $S\approx\sqrt{(n-p)/n}\approx 0.9$.
@@
Figure 4: Morphometric results for $N=51$ nominally D=30D=30\,nm UGNSs. (a) Retrieved ARs shown as filled circles with a color indicating the fitted diameter $D_{V}$. ARs retrieved using only $\sigma_{\Lambda}(\gamma_{\mathrm{P}})$ are shown as black crosses. Also shown are retrieved ARs for JC with $g=0.3$, (empty triangles), and for MP with $g=1.8$ (red circles). (b) Mean AR $\sqrt{bc}/a$ as a function of $D_{V}$, with the histogram showing the number distribution of $D_{V}$ (magenta filled circles). Data using only $\sigma_{\Lambda}$($\gamma_{\mathrm{P}}$) (cyan crosses) are also shown. The black filled squares indicate TEM measurements assuming $b=c$. Histograms are colored according to the corresponding data. Inset: pitch ($\theta$) and roll ($\phi$) angles, to show the out-of-plane orientation ($\theta\neq 0$). The size of the circles in (a–b) is given by $S^{-}=1/(1+S)$, with larger size indicating lower fit error $S$. (c) Simulated effect of the measurement noise on the retrieved ARs for individual NPs indicated by the colored stars in (a–b). Results for 1000 realisations of Gaussian noise with $\hat{\sigma}_{\mathrm{ext}}$ standard deviation added to the measured extinction data are shown. Inset : $D_{V}$ histograms, providing a colour scale for the symbols. Results for OL, JC, and MP at $g$ of $1.8$, $0.3$, and $1.8$, respectively, with corresponding results labelled. Black symbols are results using only $\sigma_{\Lambda}(\gamma_{\mathrm{P}})$ for OL.
@@
$\bar{S}$ is shown in Fig. 3 as a function of the surface damping parameter $g$ for each of the experimental $\epsilon$ datasets. We find minimum values of $\bar{S}=1.29$ at $g=1.8$ for the MP dataset, $\bar{S}=2.05$ at $g=1.8$ for the OL dataset, and $\bar{S}=1.72$ at $g=0.3$ for the JC dataset. Compared to our results in Ref. 44, $\bar{S}$ is more sensitive to $g$, due to the lower noise and larger number of measurements used. The histograms of $S$ are shown in Fig. 3 for each dataset and $g$ at the minimum of $\bar{S}(g)$ and either side of it. At the minimum, a nearly Gaussian shape is observed. The distributions of $S$ for each dataset display a few outliers of large $S$, indicating NPs are not well described by the model.
@@
Notably, the smallest $\bar{S}$ is found for MP, which indicates the best choice since it represents the model which best fits the data. On the other hand, we see in Fig. 3, LABEL:SM-fig:Sbar_SM, and LABEL:SM-fig:Ratios_SM, that OL retrieves the roundest shapes (see plots of $M\{\mathrm{\sqrt{bc}/a}\}$), closest to the shape measured in TEM. Excluding the radial polarisation data yields the lowest $\bar{S}$, for all permittivities (see Fig. LABEL:SM-fig:Sbar_SM), since the radial polarisation probes the shape in the out-of-plane direction, restraining the fit from using the out-of-plane NP shape to create spectral shifts measured for in-plane extinction.
@@
In principle, all permittivity datasets should be consistent apart from surface damping, as they have all been measured on gold films of different crystallinity. While this is the case for OL and JC, with JC corresponding to OL with $g\approx 1$, consistent with the lower crystallinity of the film measured in JC compared to the single crystal data in OL, MP corresponds to a higher plasma frequency, leading to a blueshift[44] of the LSPR of a spherical NP by about 10 nm. Using MP in the fit thus yields higher NP asymmetry, due to a compensation of that blueshift by redshift due to non-spherical shape. Going forward we thus use the OL dataset at $g=1.8$, as in our previous publication,[44] and show results using both linear and radial polarisation, unless otherwise noted.
@@
The results of the morphometric fits for the $N=51$ nominally $D=30\,\mathrm{nm}$ UGNS are shown in panels (a–b) of Fig. 4. The retrieved NP shapes, shown as filled circles, are slightly non-spherical, with $c/b=0.91\pm 0.032$, $c/a=0.82\pm 0.12$, and $b/a=0.96\pm 0.04$, where the errors are standard deviations of the ensemble. We find $\bar{S}\approx 2.1$, notably higher than $\bar{S}\approx 0.8$ found in Ref. 44, which was close to the noise-limited value. The mean AR, $\sqrt{bc}/a$, is around 0.89. When excluding $\sigma_{\Lambda}(\gamma_{\mathrm{R}})$, the results (crosses) show slightly rounder NP shapes, with $c/b=0.88\pm 0.07$, $c/a=0.85\pm 0.07$, $b/a=0.97\pm 0.03$, and a mean AR of about $0.91$. Here, the shapes form two clusters, with the smaller cluster having $c/b$ lower than any shapes retrieved when including $\sigma_{\Lambda}(\gamma_{\mathrm{R}})$. We find $D_{V}=(25.6\pm 0.5)\,$nm over the ensemble, showing a systematic deviation of about $2$ nm, or 8%, to the TEM results. The results for the JC and MP datasets, also shown in Fig. 4a, generally show retrieved shapes which are are more widely distributed and of stronger non-sphericity compared to OL.

### v2

@@ IV Results and Discussion
This work demonstrates the enhanced capabilities of the new ON including radial polarisation in determining individual NP morphometries in 3D. We maintain the naming and definitions of variables and quantities used in our previous publications,[43, 44] and introduce them in the text where needed. All analysis assumes that the refractive index of the NP environment in the samples is $n_{\mathrm{m}}=1.52$. The effect of variation of this index is discussed in Sec. F.
@@ IV.1 Statistical Distributions of $\sigma_{\Lambda}$ and $\alpha_{\Lambda}$
We first introduce relevant quantities used in the statistical description of the optical measurements. For each NP, Eq. 9 is fitted to the $\sigma_{\Lambda}(\gamma_{\mathrm{P}})$ measured for linear polarisation, determining its parameters $\sigma$, $\alpha$, and $\gamma$. Their standard errors $\hat{\sigma}$, $\hat{\alpha}$, and $\hat{\gamma}$ are determined using a Monte Carlo simulation where we add random Gaussian noise to the measured data, matching the experimentally determined noise, as detailed in Ref. 43 and Sec. C.
Given a set of $N$ NPs, numbered by the index, $i$, we identify the fit parameters using the subscript $\Lambda,i$, and define the sample-wide mean of $\sigma_{\Lambda,\,i}$ to be $\bar{\sigma}_{\Lambda}=\frac{1}{N}\sum_{i=1}^{N}\sigma_{\Lambda,\,i}$, and the associated standard deviation $\hat{\sigma}_{\Lambda,\mathrm{m}}$. Equivalent quantities are defined for the other two parameters. Correcting for the measurement noise, the standard deviation $\hat{\sigma}_{\Lambda}=\sqrt{\hat{\sigma}_{\Lambda,\mathrm{m}}^{2}-\frac{1}{N}\sum_{i=1}^{N}\hat{\sigma}_{\mathrm{\Lambda},i}^{2}}$ represents the distribution of NP cross-sections originating from their sizes and shapes only. For the relative amplitude parameter, the measurement error $\hat{\alpha}_{\Lambda,i}$ varies strongly depending on the parameter values, so that we refrain from such a correction and directly use $\hat{\alpha}_{\Lambda}=\hat{\alpha}_{\Lambda,\,\mathrm{m}}$.
@@
Fig. 2 presents $\sigma_{\Lambda}$ and $\alpha_{\Lambda}$ for measurements of the $N=51$ nominally D=30D=30\,nm UGNS for all $\Lambda$ of the experiment. To mitigate small differences in focus between the different fields of view, $\sigma_{\mathrm{ext}}$ was determined using the larger radius $r_{\mathrm{h}}$ (see Sec. D for more details). One must correct $\sigma_{\mathrm{ext}}$ obtained with a finite $r_{\mathrm{i}}$ to compensate for the signal outside of the integration area. The correction factor to $\sigma_{\mathrm{ext}}$ was determined to be $c_{\mathrm{l}}=1.12$, $c_{\mathrm{s}}=1.06$, and $c_{\mathrm{h}}=1.02$, for $r_{\mathrm{l}}$, $r_{\mathrm{s}}$, and $r_{\mathrm{h}}$, respectively. In Fig. 2 we show $\sigma_{\mathrm{ext}}$ as measured and the modelled extinction divided by $c_{\mathrm{h}}$. Narrow distributions of $\sigma_{\Lambda}$ are observed for all $\Lambda$. Small in-plane asymmetries $\alpha_{\Lambda}<0.2$ are found for about 95% of the NPs at $\Lambda={\color[rgb]{1,0.6,0}600}$ and below, and for about 80% of NPs at $\Lambda={\color[rgb]{1,0,0}650}$. At $\Lambda={\color[rgb]{0.7,0,0}700}$ instead, $\sigma_{\mathrm{ext}}$ is only a few times larger than the noise, which results in a wide distribution of $\alpha_{\Lambda}$ estimated by the shaded areas. Hence, the sample exhibits a low optical asymmetry, with some 10% outliers.
@@
Looking at the simulations in the inset of Fig. 2, we note that $\Lambda={\color[rgb]{0,0.9,0}550}$ covers the higher wavelength side of the LSPR of a spherical gold NP, suggesting that accurately modelling the width of the channels will be relevant for the morphology retrieval. With increasing NP asymmetry, the LSPR polarised along the longer axis (the longitudinal mode) red-shifts, increasing its overlap with the $\color[rgb]{0,0.9,0}550$} channel, and thus providing a sensitive probe of asymmetries. We find $\bar{\alpha}_{{\color[rgb]{0,0.9,0}550}}=0.037\,\pm\,0.039$, with some 80% of NPs showing $\alpha_{{\color[rgb]{0,0.9,0}550}}<0.054$, and the rest clearly asymmetric with $\alpha_{{\color[rgb]{0,0.9,0}550}}$ up to 0.194. For asymmetric NPs, the LSPR will move into the increasingly red channels (see simulations in Ref. 43 Fig. 2), increasing their $\sigma_{\Lambda}$ and $\alpha_{\Lambda}$. Looking at the exemplary TEM image, some 10-20% of significantly non-spherical NPs can be expected. Their red-shift can be significant, contributing to the values of $\alpha_{{\color[rgb]{1,0,0}650}}$ and $\alpha_{{\color[rgb]{0.7,0,0}700}}$ well above the noise. We also note that the $\Lambda=\{{\color[rgb]{0,0,1}450},\,{\color[rgb]{0,0.92,0.92}500}\}$ channels are rather insensitive to asymmetry, due to the dominant absorptive part of the permittivity, $\Im\{\epsilon_{\mathrm{NP}}\}\approx 5$, with $\Re\{\epsilon_{\mathrm{NP}}\}$ similar to $\epsilon_{\mathrm{m}}$, suppressing shape-dependent effects, so that $\sigma_{{\color[rgb]{0,0,1}450}}$ and $\sigma_{{\color[rgb]{0,0.92,0.92}500}}$ are good reporters of the NP volume.
@@
Let us now discuss $\sigma_{\mathrm{ext}}$ measured using the radial polariser, $\sigma_{\Lambda}(\gamma_{\mathrm{R}})$, see vertical bars in Fig. 2, which we compare to the fitted $\sigma_{\Lambda}$ for linear polarisation shown as crosses. We recall that the electric field at the sample has both in-plane and out-of-plane components, and we consider[35, 52] the relative strengths of the components along the $x$, $y$, and $z$ axes, respectively as $E=[E_{\mathrm{x}},\,E_{\mathrm{y}},\,E_{\mathrm{z}}]$. The intensity in the back focal plane (BFP) of the condenser varies spatially, with a maximum at the center, determined by the lamp, diffuser and lenses in the illumination path, as characterised in our previous work.[35] Using the known spatial dependence, we can calculate the intensity components at the sample polarised along the $x$, $y$, and $z$ axes, normalized to have unity sum. For a linear polariser along $x$ in the condenser BFP, we find $I_{\mathrm{lin}}=[0.826,\,0.007,\,0.167]$, corresponding to field components $E_{\mathrm{lin}}=[0.909,\,0.084,\,0.409]$ given by the square root of the intensity components. For radially polarised input to the BFP, we have $I_{\mathrm{rad}}=[0.333,0.333,0.334]$ corresponding to $E_{\mathrm{rad}}=[0.577,\,0.577,\,0.578]$, showing that the $z$ component is increased in the radially polarised case.
@@
For a spherical NP, the polarisation direction should not affect the measured $\sigma_{\mathrm{ext}}$, and thus we expect the difference $\delta_{\sigma}=\sigma_{\Lambda}(\gamma_{\mathrm{R}})-\sigma_{\Lambda}$ to be negligible. The histograms of $\delta_{\sigma}$ are shown in Fig. 2b, and we find for channels with large cross-sections $\Lambda=\{{\color[rgb]{0,0,1}450},\,{\color[rgb]{0,0.92,0.92}500},\,{\color[rgb]{0,0.9,0}550}\}$ a shift towards $\delta_{\sigma}>0$ of relative magnitude $\delta_{\sigma}/\sigma_{\Lambda}$ below 3%, while for the other channels $\delta_{\sigma}$ is dominated by noise, shown as gray bars. To explain this small shift, we recall the “long shadow” effect,[55] which arises from the illumination covering an oblique angular range while referencing to the intensity transmitted through the imaged sample plane. The $\mathrm{NA}$ of the condenser and the intensity distribution in its BFP provide the weighting of the illumination angles at the sample, from which we calculate the long-shadow factor used. However, the transmission of the condenser and objective for oblique angles is polarisation-dependent, due to the reflections on the lens surfaces, which are generally lower for p polarisation compared to s polarisation. Radial polarisation is purely p polarised, while linear polarisation contains equal parts of s and p polarised light. Considering that the long-shadow factor is increasing with incidence angle,[55] the higher transmission at higher angles for radial polarisation creates an illumination with a higher long-shadow factor. This systematic error can be approximately compensated by reducing the measured $\sigma_{\Lambda}(\gamma_{\mathrm{R}})$ by a factor $0.97$, having an effect explored below. One could also speculate that a preferential orientation of the NPs on the surface creates the difference. However, we would expect in such a case that NPs preferentially lie flat on the surface, providing a larger binding surface, which would lead to a higher cross-section for in-plane fields, opposite to what is observed.
@@
Figure 3: $\bar{S}$ and selected statistics of $S$ for $N=51$ UGNS of nominal D=30D=30\,nm for varying surface damping parameter $g$ and different $\epsilon$ datasets as indicated by color. (a) Top: $\bar{S}$ versus $g$. Solid lines and squares for $\epsilon_{\mathrm{NP}}(\lambda)$. Dotted lines and stars for $\epsilon_{\mathrm{NP}}^{s}(\lambda)$ with $0.97\times$ correction for $\sigma_{\mathrm{ext}}(\gamma_{\mathrm{R}})$ data. Bottom: median of the fitted average aspect ratio, for MAR (filled symbols) and PAR (empty symbols). Solid and dotted line designations same as Top. (b–j) Histograms of $S$, with (b–d) using the OL dataset with $g=\{0,\,1.8,\,3.6\}$, (e–g) JC dataset with $g=\{0,\,0.3,\,1.8\}$, (h–j) MP dataset with $g=\{0,\,1.8,\,3.6\}$. The values of $g$ shown are at $g=0$, the minimum of $\bar{S}(g)$ and a larger $g$.
@@ IV.2 Morphometric analysis
where we dropped the index $i$ for brevity. $n$ is the number of measurements in the sum, which includes the radially polarised measurement $\gamma_{\mathrm{P}}=\gamma_{\mathrm{R}}$, and $\hat{\sigma}_{\Lambda}(\gamma_{\mathrm{P}})$ is the measurement noise associated with $\sigma_{\Lambda}$($\gamma_{\mathrm{P}}$). The factor $\eta$ is the fraction of scattered light which is not collected by the objective and thus contributes to the measured extinction. For our experimental configuration we calculated [54] $\eta=0.646$.
@@
Since $S$ is normalized by $\hat{\sigma}_{\mathrm{ext}}(\gamma_{\mathrm{P}})$ in Eq. 11, we would expect $S=1$ for the correct model, without taking into account the number of fit parameters and the number of data points to be fitted per NP. Considering that we use $p=6$ free parameters to fit the $n=30$ measurements (four linear and one radial polarisation at 6 wavelengths), this expectation is reduced by the number of degrees of freedom npn-p to $S\approx\sqrt{(n-p)/n}\approx 0.9$.
@@
Figure 4: Morphometric results for $N=51$ nominally D=30D=30\,nm UGNSs. (a) Retrieved ARs for OL $g=1.8$ are shown as filled symbols with a colour indicating the fitted diameter $D_{V}$. ARs retrieved using only $\sigma_{\Lambda}(\gamma_{\mathrm{P}})$ are shown as cyan empty diamonds. Also shown are retrieved ARs for JC $g=0.3$, (magenta empty triangles), and for MP $g=1.8$ (red empty circles). (b) MAR versus $D_{V}$ for OL $g=1.8$ (magenta empty circles), with the histogram showing the number distribution of $D_{V}$ (empty) and $D_{\mathrm{p}}$ (filled). Corresponding PAR are shown as filled symbols versus $D_{\mathrm{p}}$. Results using only $\sigma_{\Lambda}$($\gamma_{\mathrm{P}}$) are shown as diamonds. The black filled squares indicate TEM measurements PAR versus $D_{\mathrm{p}}$. Inset: pitch ($\theta$) and roll ($\phi$) angles showing the particle orientation for OL, with a symbol colour encoding the MAR. The size of the circles in (a) and the empty symbols in (b) is given by $S^{-}=1/(1+S)$, as indicated in (a), with larger size indicating lower fit error $S$. (c) Simulated effect of the measurement noise on the retrieved ARs for individual NPs indicated by the colored stars in (a). Results for 1000 realisations of Gaussian noise with $\hat{\sigma}_{\mathrm{ext}}$ standard deviation added to the measured extinction data are shown. Inset : $D_{V}$ histograms, providing a colour scale for the symbols. Results for OL, JC, and MP at $g$ of $1.8$, $0.3$, and $1.8$, respectively, with corresponding results labelled. Black symbols are results using only $\sigma_{\Lambda}(\gamma_{\mathrm{P}})$ for OL.
@@
$\bar{S}$ is shown for the data analysis as discussed to this point by the solid curves in Fig. 3a as a function of the surface damping parameter $g$ for each of the experimental $\epsilon$ datasets. We find minimum values of $\bar{S}=1.29$ at $g=1.8$ for the MP dataset, $\bar{S}=2.05$ at $g=1.8$ for the OL dataset, and $\bar{S}=1.72$ at $g=0.3$ for the JC dataset. Compared to our results in Ref. 44, $\bar{S}$ is more sensitive to $g$, due to the lower noise and larger number of measurements used. The histograms of $S$ are shown in Fig. 3 for each dataset and $g$ at the minimum of $\bar{S}(g)$ and either side of it. At the minimum, a nearly Gaussian shape is observed. The distributions of $S$ for each dataset display a few outliers of large $S$, indicating some NPs are not well described by the model.
@@
Notably, the smallest $\bar{S}$ is found for MP, which indicates the best choice since it represents the model which best fits the data. On the other hand, we see in Fig. 3, 12, and 13, that OL retrieves the roundest shapes as shown by $M\{\mathrm{MAR}\}$, closest to the shape measured in TEM. Excluding the radial polarisation data yields the lowest $\bar{S}$, for all permittivities (see Fig. 12). This can be rationalised noting that the radial polarisation probes the shape in the out-of-plane direction, restraining the fit from using this shape to create spectral shifts while the in-plane shape is constrained by $\sigma_{\Lambda}(\gamma_{\mathrm{P}})$.
@@
In principle, all permittivity datasets should be consistent apart from surface damping, as they have all been measured on gold films of different crystallinity. While this is the case for OL and JC, with JC corresponding to OL with $g\approx 1$, reflecting the lower crystallinity of the film measured in JC compared to the single crystal data in OL, MP corresponds to a higher plasma frequency, leading to a blueshift of the LSPR of a spherical NP by about[44] 4 nm. Using MP in the fit thus yields higher NP asymmetry, which compensates the blueshift by a redshift due to non-spherical shape.
@@
The results of the morphometric fits for the $N=51$ nominally $D=30\,\mathrm{nm}$ UGNS are shown in panels (a–b) of Fig. 4. The retrieved NP shapes for OL $g=1.8$, shown as filled circles, are slightly non-spherical, with $c/b=0.91\pm 0.032$, $c/a=0.82\pm 0.12$, and $b/a=0.96\pm 0.04$, where the errors are standard deviations of the ensemble, and a mean (median) MAR of 0.89 (0.90). When excluding $\sigma_{\Lambda}(\gamma_{\mathrm{R}})$, the results show slightly rounder NP shapes, with $c/b=0.88\pm 0.07$, $c/a=0.85\pm 0.07$, $b/a=0.97\pm 0.03$, and a mean (median) MAR of 0.91 (0.91). Here, the shapes form two clusters, with the smaller cluster having $c/b$ lower than any shapes retrieved when including $\sigma_{\Lambda}(\gamma_{\mathrm{R}})$. We find $D_{V}=(25.6\pm 0.5)\,$nm (median 25.6 nm) over the ensemble when including $\sigma_{\Lambda}(\gamma_{\mathrm{R}})$, showing a systematic deviation of about $2$ nm, or 8%, to the TEM results. The results for the JC and MP datasets generally show retrieved shapes which are more widely distributed and of stronger non-sphericity compared to OL, and MP retrieves nearly $1\,\mathrm{nm}$ larger average diameters. We explore in Sec. E the effect of the different corrections and selections introduced on the shape retrieval in more detail.
Notably, the projected shapes are generally rounder and of larger diameter, see filled circles (PAR) and filled histogram ($D_{\mathrm{p}}$) in Fig. 4b. This can be interpreted in two ways: (i) since the in-plane asymmetry is measured with higher precision via the $\gamma_{\mathrm{P}}$ dependence, the PAR has less error and thus reflects the correct aspect ratio in all three dimensions of the NP, (ii) the NPs are orientated on the surface in a way to appear more round in-plane. Considering the nearly spherical shapes, a physical mechanism which would result in (ii) both for TEM and optical measurements is not obvious.